# Feasibility Study on Active Learning of Smart Surrogates for Scientific Simulations

### v1

@@ Abstract.
Active Learning, Diffusion Solver, Deep Learning, Surrogate Modeling
copyright: rightsretainedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Platform for Advanced Scientific Computing; Zurich, Switzerland; submissionid: ccs: Computing methodologies Active learning settingsccs: Computing methodologies Machine learning
1. Introduction
@@
However, despite HPC advances, a limitation arises as the computational demands of the increasingly complex simulation models surpass the system’s capacity, especially when exploring extensive parameter spaces. Recent developments in machine learning and deep learning have sparked an interest in creating efficient surrogate approximations for these scientific simulations. For either certain internal components of the simulations or the entire simulations, these surrogates could drastically accelerate expensive scientific simulations by several orders of magnitude (Fox and Jha, 2019; Jha and Fox, 2019; Fox et al., 2019). Amidst these, deep neural network (DNN) based surrogates have emerged as powerful tools for optimizing high-performance scientific simulations in multidimensional space (Kasim et al., 2021; Hruska et al., 2020; Lee et al., 2019; Mardt et al., 2018; Bhowmik et al., 2018; Ton et al., 2020; Liao et al., 2019; Clyde et al., 2020; Sun et al., 2020; Gao et al., 2021; Mustafa et al., 2019). These approaches address both the computational challenges associated with extensive simulations and help improve the efficiency and speed of scientific inquiry.
@@
In this paper, we attempt to provide initial insights into the above critical gaps in the specific context of diffusion equations with sources. Often in science, we are interested in the rate of change of some physical quantity concerning time and how this rate of change depends on the environment. The most common way to model the process is via partial differential equations. This concept underpins a myriad of phenomena, exemplified by models such as the diffusion equation in heat conduction, Schrödinger’s equation in quantum mechanics, the Navier-Stokes equation in fluid dynamics, and others in fields ranging from electromagnetism to finance and biology. The intricacies of these problems are often dictated by their nature and specific initial and boundary conditions, with solutions varying from simple stationary answers to complex dynamic responses. Here we focus on one such instance as a use case – the stationary solution of the diffusion equation in scenarios with sources. Despite their ubiquity across various systems, diffusion processes pose substantial computational challenges, particularly when solving steady-state and time-varying equations in environments with parameters that vary significantly. These complexities are compounded when factors like the number of sources and sinks fluctuate, increasing the computational demands of simulations (Toledo-Marín et al., 2021). DNNs have been proposed as PDE solvers (van den Dool et al., 2023). In particular, convolutional neural networks have been successful mainly due to the inherent position encoding in convolutional layers. Alterantively, U-Nets that were first proposed for image segmentation (Ronneberger et al., 2015) have been used in as PDE solvers as well (Gupta and Brandstetter, 2022). We build on these past successes and study the effectiveness of different active learning strategies in building DNN surrogates for this problem, considering both the CNN and U-Net architectures reported in previous works.
@@
2. Related Works
Diffusion Solvers: Diffusion solvers are integral to scientific computing, given the widespread occurrence of diffusion processes across various domains. Traditional numerical methods, such as Finite Elements and Finite Differences, have been the cornerstone in this area due to their robustness and efficiency in certain scenarios (Belytschko, 2007; Schiesser, 2012) and, in some cases, these methods can be fairly optimized (Rackauckas and Nie, 2017). However, the effectiveness of these methods often hinges on specific factors such as the geometry of the problem, inherent symmetries, and the problem parameters (e.g., diffusivity field, decay rate field, etc.), which can limit their applicability in more complex scenarios. In response to these limitations, DNNs have emerged as promising surrogates, offering the potential to expedite simulations traditionally handled by these numerical methods (Farimani et al., 2017; Han et al., 2018; Li et al., 2020; He and Pathak, 2020; Willard et al., 2020; Cai et al., 2021). Despite their promise, deploying neural networks in this context is challenging. Issues such as prolonged training times, the scarcity of curated datasets, and a lack of physical principles integration can significantly impede their performance and reliability in accurately modeling diffusion processes (Toledo-Marin et al., 2023). Addressing these challenges is crucial for effectively integrating neural networks in diffusion solvers, ensuring both speed and accuracy in scientific computations.
@@ 2. Related Works
Deep Active Learning (DAL): Active learning is an area in machine learning that deals with incremental intelligent selection of data to query for labels to achieve high model performance with low annotation cost (Cohn et al., 1996). In the context of deep learning, this practice is referred to as Deep Active Learning (DAL), which has led to the development of various data selection strategies. The strategies can be categorized into uncertainty-based, diversity-based, and hybrid methods. Uncertainty-based methods actively select examples that the DNN is most uncertain about. Examples of such methods include estimation of DNN loss (Yoo and Kweon, 2019; Huang et al., 2021), entropy (Joshi et al., 2009; Gal and Ghahramani, 2016), BALD (Houlsby et al., 2011; Kirsch et al., 2019), margin sampling (Roth and Small, 2006; Scheffer et al., 2001), and least confidence sampling based on softmax output (Settles, 2009). The diversity-based method aim to find a subset of diverse samples that effectively represents the data distribution. These methods often rely on either core set selection (Sener and Savarese, 2017; Geifman and El-Yaniv, 2017) or density-based clustering (Nguyen and Smeulders, 2004; Wang et al., 2017) to find examples that are most representative of the distribution of the entire unlabeled pool of data. Hybrid methods combine the benefit of both uncertainty-based and diversity-based approaches (Wang et al., 2016; Zhou et al., 2017; Yang et al., 2017; Ash et al., 2019; Shui et al., 2020; Wang et al., 2022; Kong et al., 2022).
@@
Active Learning in Surrogate Modeling: Some recent works have emerged in exploring the role of active learning in DNN surrogates for scientific simulations. A Bayesian active learning was proposed in (Wu et al., 2023) to use the expected information gain to select training simulations when training a neural process as the surrogate model of scientific simulations, tested on the reaction-diffusion, heat, and local epidemic and mobility models. An active learning method is presented in (Pestourie et al., 2020) to reduce the number of simulations for an NN surrogate model of optical-surface components for photonic device modeling. They suggests an ensemble training-based acquisition function which selects the examples based on the error measure. In (Lye et al., 2021), a DNN surrogate is built to link the PDE parameter to the observables in the context of solving for PDE constrained optimizations, where the acquisition function is based on the optimization objective function using the DNN surrogates and selects ’k’ parameters below the tolerance set. A two network (classifier and regressor) physics-informed active learning system, named ADEPT (Active Deep Ensembles for Plasma Turbulence), is presented in (Zanisi et al., 2023). This framework is designed to train a surrogate model for gyrokinetic turbulence, with the primary goal of minimizing the required data volume. The classifier screen identifies potential candidates from unlabelled pool and regressor’s epistemic uncertainty is used to strategically select the examples for simulation.
@@
3. Background: Diffusion Solver
@@
Figure 1. Distribution of parameters used to generate simulation data. (a) Histogram of distance between two sources on the lattice. (b) Histogram of intensity of source 2 which has randomly assigned value between 0 and 1. (c-d): Histograms of the positions of (c) source 1 and (d) source 2 on the lattice.
3.1. Data Generation
The data used in this paper is motivated by biological multiscale modeling (Toledo-Marín et al., 2021), and correspond to two sources placed in a lattice with a decay rate which acts as a sink. To generate representative initial conditions and corresponding steady-state diffusion fields from a two-source system, we consider a 100x100 $unit^{2}$ lattice. In total, 20k initial configurations were created with two sources randomly positioned in the lattice. Each source, characterized by a 5-unit radius, has a constant flux of 1 for one source, while the other sources have a randomly assigned constant flux value between 0 and 1, randomly assigned using a uniform distribution. The remainder of the lattice has a field value of zero. The stationary solution to the diffusion equation with absorbing boundary conditions for each initial condition is computed using the Differential Equation package in Julia (Rackauckas and Nie, 2017). We denote $x$ as the input image that represents the initial condition layout of the two source cells, and $y$ represents the output image, which is the predicted stationary solution of the diffusion equation.
@@ 3.1. Data Generation
3.2. DNN Surrogates
We define two DNN architectures, namely, a CNN based autoencoder and an U-Net-based structure as shown in Figs.  2 and  3, respectively.
The autoencoder’s first six layers perform an average-pool operation that reduces height and width in half after each layer following the sequence {$100^{2}$, $50^{2}$, $25^{2}$, $12^{2}$, $6^{2}$, $3^{2}$, $1^{2}$} while adding channels after each layer following the sequence {1, 64, 128, 256, 512, 1024, 2048}. Similar to (Toledo-Marín et al., 2021), the next six layers reduce the number of channels while increasing the height and width in the opposite order from those described above.
@@ 3.2. DNN Surrogates
Figure 2. Deep convolutional neural network architecture for the diffusion surrogate. All convolution layers leave the input with the same height and width. Each block is composed of a convolution which increases the channels from $Ch1$ to $Ch2$, followed by a LeakyReLU with slope set at $0.02$.
Figure 3. U-Net architecture for the diffusion surrogate. All convolution layers leave the input with the same height and width. Each block comprises a convolution that increases the channels from $Ch1$ to $Ch2$, followed by a ReLU, a convolution that leaves the dimensions unchanged, and a ReLU afterward. The main feature in U-Nets is the concatenation between blocks in the encoder and in the decoder.
3.3. Loss Definition
@@
(1) $L^{\alpha}_{i\beta}=\exp\left(-\frac{(\langle i|1\rangle-\langle i|y_{\beta}\rangle)}{w}\right)\cdot(\langle i|\hat{y}_{\beta}\rangle-\langle i|y_{\beta}\rangle)^{\alpha}$
@@
4. Methods: Active Learning of DNN Surrogates for Diffusion Solvers
4.1. Active Learning
@@
Note that an important consideration for the design of acquisition functions is that the labels (in this case the simulation outputs on the given inputs) of the samples are not available before the samples are selected to be queried for label. In another word, the acquisition objective cannot utilize the label of the samples. We consider three active learning acquisition functions and compare them with random acquisition. Two are representative of uncertainty-based acquisitions, including the use of entropy (Joshi et al., 2009) and the estimated loss of the DNN surrogage (Huang et al., 2021). One is representative of acquisitions considering the diversity of training samples.
@@
Entropy sampling (Joshi et al., 2009) selects instances that the model is most uncertain about as measured by the standard deviation of the ’k’ output predictions of the network. We use dropout as Bayesian approximation (Gal and Ghahramani, 2016) to generate ’k’ different output predictions and evaluate entropy as:
(2) $Entropy=\frac{1}{Nk}\sum_{j=1}^{N}\sum_{r\in\{1,2,...,k\}}(\hat{y}_{rj}-\bar{y}_{j})^{2}$
@@
The loss of the DNN surrogate on a new sample is another potential criteria for selecting samples a DNN is most uncertain about: a high-loss sample has the potential to provide stronger signals for the optimization of the DNN, whereas a low-loss sample could be redundant to what the DNN is already trained with. Intended as a criteria for selecting unlabeled data samples, the loss of the DNN on the unlabeled samples cannot be calculated from the actual labels (here the simulation outputs) but must be estimated. Following the Temporal Output Discrepancy (TOD) measure (Huang et al., 2021), we estimate the DNN loss on a sample based on the discrepancy of the DNN outputs on the same sample at different learning iterations:
(3) $D_{t}^{T}(x)\overset{\mathrm{def}}{=}||f(x;\theta_{t+T}-f(x;\theta_{t})||$
@@
Figure 4. Regions of interest shown with white for a given lattice – a) Source, b) Field, c) Ring 1, d) Ring 2 and e) Ring 3 for two sources placed randomly in the lattice.
Figure 5. Comparison of test weighted-MAE on a) U-Net architecture and b) CNN autoencoder, in log scale, between the ground truth and prediction from the DNN surrogates across different acquisition functions as data acquisition proceeds.
@@
To consider the diversity of training simulations without looking at the simulated outputs, we measure diversity by the parameters used to generate the simulations. We consider six parameters in the input initial condition, including the coordinates (cx and cy) of the two sources ($cx_{1},cy_{1},cx_{2},cy_{2}$), the distance between the sources ($d$), and the intensity of the second source ($Q$) as the important identifiers ($I$) for the model. Following (Sener and Savarese, 2017), we create a coreset subset from the unlabeled data such that the examples with the largest minimal distance from the labeled data are selected.
(4) $dist=max_{m\in U}\ min_{n\in L}\bigtriangleup(I_{m},I_{n})$
@@
4.2. Evaluation Metrics
Figure 6. Comparison of test weighted-MAE on a) U-Net architecture and b) CNN autoencoder, in log scale, between the ground truth and prediction from the DNN surrogates between data acquisitions using TOD and labelled-loss as data acquisition proceeds.
@@ 4.2. Evaluation Metrics
As mentioned earlier, the sources occupy 2% of the total lattice space where the diffusion pattern emanate from the source and die out quickly in the field. To better understand the effect of active learning on the accuracy of DNN surrogates, we further examine the MAE in different regions of interest on the lattice as illustrated in Figure  4: the regions of ”Source” and ”Field” are determined based on the initial inputs to the simulations; the regions of ”RINGS 1,” ”RINGS 2,” and ”RINGS 3” are defined based on the ground-truth simulation outputs using pixel values in the ranges [0.2, 1.0], [0.1, 0.2], and [0.05, 0.1], respectively – representing increasing distances to the sources.

### v2

@@ Abstract
Keywords Machine Learning  $\cdot$ Deep Learning  $\cdot$ Active Learning  $\cdot$ Diffusion Solver  $\cdot$ Surrogate Modeling  $\cdot$ Encoder-decoder
1 Introduction
@@
However, despite HPC advances, a limitation arises as the computational demands of the increasingly complex simulation models surpass the system’s capacity, especially when exploring extensive parameter spaces. Recent developments in machine learning and deep learning have sparked an interest in creating efficient surrogate approximations for these scientific simulations. For either certain internal components of the simulations or the entire simulations, these surrogates could drastically accelerate expensive scientific simulations by several orders of magnitude [1, 2, 3]. Amidst these, deep neural network (DNN) based surrogates have emerged as powerful tools for optimizing high-performance scientific simulations in multidimensional space [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. These approaches address both the computational challenges associated with extensive simulations and help improve the efficiency and speed of scientific inquiry.
@@
In this paper, we attempt to provide initial insights into the above critical gaps in the specific context of diffusion equations with sources. Often in science, we are interested in the rate of change of some physical quantity concerning time and how this rate of change depends on the environment. The most common way to model the process is via partial differential equations. This concept underpins a myriad of phenomena, exemplified by models such as the diffusion equation in heat conduction, Schrödinger’s equation in quantum mechanics, the Navier-Stokes equation in fluid dynamics, and others in fields ranging from electromagnetism to finance and biology. The intricacies of these problems are often dictated by their nature and specific initial and boundary conditions, with solutions varying from simple stationary answers to complex dynamic responses. Here we focus on one such instance as a use case – the stationary solution of the diffusion equation in scenarios with sources. Despite their ubiquity across various systems, diffusion processes pose substantial computational challenges, particularly when solving steady-state and time-varying equations in environments with parameters that vary significantly. These complexities are compounded when factors like the number of sources and sinks fluctuate, increasing the computational demands of simulations [15]. DNNs have been proposed as PDE solvers [16]. In particular, convolutional neural networks have been successful mainly due to the inherent position encoding in convolutional layers. Alterantively, U-Nets that were first proposed for image segmentation [17] have been used in as PDE solvers as well [18]. We build on these past successes and study the effectiveness of different active learning strategies in building DNN surrogates for this problem, considering both the CNN and U-Net architectures reported in previous works.
@@
2 Related Works
Diffusion Solvers: Diffusion solvers are integral to scientific computing, given the widespread occurrence of diffusion processes across various domains. Traditional numerical methods, such as Finite Elements and Finite Differences, have been the cornerstone in this area due to their robustness and efficiency in certain scenarios [19, 20] and, in some cases, these methods can be fairly optimized [21]. However, the effectiveness of these methods often hinges on specific factors such as the geometry of the problem, inherent symmetries, and the problem parameters (e.g., diffusivity field, decay rate field, etc.), which can limit their applicability in more complex scenarios. In response to these limitations, DNNs have emerged as promising surrogates, offering the potential to expedite simulations traditionally handled by these numerical methods [22, 23, 24, 25, 26, 27]. Despite their promise, deploying neural networks in this context is challenging. Issues such as prolonged training times, the scarcity of curated datasets, and a lack of physical principles integration can significantly impede their performance and reliability in accurately modeling diffusion processes [28]. Addressing these challenges is crucial for effectively integrating neural networks in diffusion solvers, ensuring both speed and accuracy in scientific computations.
@@ 2 Related Works
Deep Active Learning (DAL): Active learning is an area in machine learning that deals with incremental intelligent selection of data to query for labels to achieve high model performance with low annotation cost [29]. In the context of deep learning, this practice is referred to as Deep Active Learning (DAL), which has led to the development of various data selection strategies. The strategies can be categorized into uncertainty-based, diversity-based, and hybrid methods. Uncertainty-based methods actively select examples that the DNN is most uncertain about. Examples of such methods include estimation of DNN loss [30, 31], entropy [32, 33], BALD [34, 35], margin sampling [36, 37], and least confidence sampling based on softmax output [38]. The diversity-based method aim to find a subset of diverse samples that effectively represents the data distribution. These methods often rely on either core set selection [39, 40] or density-based clustering [41, 42] to find examples that are most representative of the distribution of the entire unlabeled pool of data. Hybrid methods combine the benefit of both uncertainty-based and diversity-based approaches [43, 44, 45, 46, 47, 48, 49].
@@
Active Learning in Surrogate Modeling: Some recent works have emerged in exploring the role of active learning in DNN surrogates for scientific simulations. A Bayesian active learning was proposed in [50] to use the expected information gain to select training simulations when training a neural process as the surrogate model of scientific simulations, tested on the reaction-diffusion, heat, and local epidemic and mobility models. An active learning method is presented in [51] to reduce the number of simulations for an NN surrogate model of optical-surface components for photonic device modeling. They suggests an ensemble training-based acquisition function which selects the examples based on the error measure. In [52], a DNN surrogate is built to link the PDE parameter to the observables in the context of solving for PDE constrained optimizations, where the acquisition function is based on the optimization objective function using the DNN surrogates and selects ’k’ parameters below the tolerance set. A two network (classifier and regressor) physics-informed active learning system, named ADEPT (Active Deep Ensembles for Plasma Turbulence), is presented in [53]. This framework is designed to train a surrogate model for gyrokinetic turbulence, with the primary goal of minimizing the required data volume. The classifier screen identifies potential candidates from unlabelled pool and regressor’s epistemic uncertainty is used to strategically select the examples for simulation.
@@
3 Background: Diffusion Solver
@@
Figure 1: Distribution of parameters used to generate simulation data. (a) Histogram of distance between two sources on the lattice. (b) Histogram of intensity of source 2 which has randomly assigned value between 0 and 1. (c-d): Histograms of the positions of (c) source 1 and (d) source 2 on the lattice.
Figure 2: Deep convolutional neural network architecture for the diffusion surrogate. All convolution layers leave the input with the same height and width. Each block is composed of a convolution which increases the channels from $Ch1$ to $Ch2$, followed by a LeakyReLU with slope set at $0.02$.
Figure 3: U-Net architecture for the diffusion surrogate. All convolution layers leave the input with the same height and width. Each block comprises a convolution that increases the channels from $Ch1$ to $Ch2$, followed by a ReLU, a convolution that leaves the dimensions unchanged, and a ReLU afterward. The main feature in U-Nets is the concatenation between blocks in the encoder and in the decoder.
3.1 Data Generation
The data used in this paper is motivated by biological multiscale modeling [15], and correspond to two sources placed in a lattice with a decay rate which acts as a sink. To generate representative initial conditions and corresponding steady-state diffusion fields from a two-source system, we consider a 100x100 $unit^{2}$ lattice. In total, 20k initial configurations were created with two sources randomly positioned in the lattice. Each source, characterized by a 5-unit radius, has a constant flux of 1 for one source, while the other sources have a randomly assigned constant flux value between 0 and 1, randomly assigned using a uniform distribution. The remainder of the lattice has a field value of zero. The stationary solution to the diffusion equation with absorbing boundary conditions for each initial condition is computed using the Differential Equation package in Julia [21]. We denote $x$ as the input image that represents the initial condition layout of the two source cells, and $y$ represents the output image, which is the predicted stationary solution of the diffusion equation.
@@ 3.1 Data Generation
3.2 DNN Surrogates
We define two DNN architectures, namely, a CNN based autoencoder and an U-Net-based structure as shown in Figure 3 and  3, respectively.
The autoencoder’s first six layers perform an average-pool operation that reduces height and width in half after each layer following the sequence {$100^{2}$, $50^{2}$, $25^{2}$, $12^{2}$, $6^{2}$, $3^{2}$, $1^{2}$} while adding channels after each layer following the sequence {1, 64, 128, 256, 512, 1024, 2048}. Similar to [15], the next six layers reduce the number of channels while increasing the height and width in the opposite order from those described above.
@@ 3.2 DNN Surrogates
3.3 Loss Definition
@@
$L^{\alpha}_{i\beta}=\exp\left(-\frac{(\langle i|1\rangle-\langle i|y_{\beta}\rangle)}{w}\right)\cdot(\langle i|\hat{y}_{\beta}\rangle-\langle i|y_{\beta}\rangle)^{\alpha}$ (1)
@@
4 Methods: Active Learning of DNN Surrogates for Diffusion Solvers
4.1 Active Learning
@@
Note that an important consideration for the design of acquisition functions is that the labels (in this case the simulation outputs on the given inputs) of the samples are not available before the samples are selected to be queried for label. In another word, the acquisition objective cannot utilize the label of the samples. We consider three active learning acquisition functions and compare them with random acquisition. Two are representative of uncertainty-based acquisitions, including the use of entropy [32] and the estimated loss of the DNN surrogage [31]. One is representative of acquisitions considering the diversity of training samples.
@@
Entropy sampling [32] selects instances that the model is most uncertain about as measured by the standard deviation of the ’k’ output predictions of the network. We use dropout as Bayesian approximation [33] to generate ’k’ different output predictions and evaluate entropy as:
$Entropy=\frac{1}{Nk}\sum_{j=1}^{N}\sum_{r\in\{1,2,...,k\}}(\hat{y}_{rj}-\bar{y}_{j})^{2}$ (2)
@@
Figure 4: Regions of interest shown with white for a given lattice – a) Source, b) Field, c) Ring 1, d) Ring 2 and e) Ring 3 for two sources placed randomly in the lattice.
@@
The loss of the DNN surrogate on a new sample is another potential criteria for selecting samples a DNN is most uncertain about: a high-loss sample has the potential to provide stronger signals for the optimization of the DNN, whereas a low-loss sample could be redundant to what the DNN is already trained with. Intended as a criteria for selecting unlabeled data samples, the loss of the DNN on the unlabeled samples cannot be calculated from the actual labels (here the simulation outputs) but must be estimated. Following the Temporal Output Discrepancy (TOD) measure [31], we estimate the DNN loss on a sample based on the discrepancy of the DNN outputs on the same sample at different learning iterations:
$D_{t}^{T}(x)\overset{\mathrm{def}}{=}||f(x;\theta_{t+T}-f(x;\theta_{t})||$ (3)
@@
To consider the diversity of training simulations without looking at the simulated outputs, we measure diversity by the parameters used to generate the simulations. We consider six parameters in the input initial condition, including the coordinates (cx and cy) of the two sources ($cx_{1},cy_{1},cx_{2},cy_{2}$), the distance between the sources ($d$), and the intensity of the second source ($Q$) as the important identifiers ($I$) for the model. Following [39], we create a coreset subset from the unlabeled data such that the examples with the largest minimal distance from the labeled data are selected.
$dist=max_{m\in U}\ min_{n\in L}\bigtriangleup(I_{m},I_{n})$ (4)
@@
4.2 Evaluation Metrics
@@
As mentioned earlier, the sources occupy 2% of the total lattice space where the diffusion pattern emanate from the source and die out quickly in the field. To better understand the effect of active learning on the accuracy of DNN surrogates, we further examine the MAE in different regions of interest on the lattice as illustrated in Figure 4: the regions of "Source" and "Field" are determined based on the initial inputs to the simulations; the regions of "RINGS 1," "RINGS 2," and "RINGS 3" are defined based on the ground-truth simulation outputs using pixel values in the ranges [0.2, 1.0], [0.1, 0.2], and [0.05, 0.1], respectively – representing increasing distances to the sources.
5 Experiments
We split the two-source data described in section 3.1. into 16,000 training data, 4,000 validation data, and 4,000 test data. We trained the U-Net and CNN autoencoder structure defined in section 3.2. We divided the 16000 training data into an initial 1000 labelled data (samples with simulation outputs available) and a 15000 unlabelled pool (samples without simulation outputs). We performed active learning with the acquisition functions defined in section 4.1 for CNN autoencoder and exclude entropy acquisition function for U-Net due to large computational demand for ${}^{\prime}N^{\prime}$ forward passes. At each acquisition round, new unlabled samples are selected to query for simulation outputs, and the DNN surrogates were retrained with the current labeled dataset (input-output simulationp pairs). This was repeated until all the 15000 unlabelled samples are labeled. In each active learning round, the DNN surrogate was trained for 500 epochs and saved based on the best validation loss. For the entropy acquisition function, in CNN autoencoder, 40% dropouts were added after batch normalization of the first and second convolutional layers. The code repository can be found in [56].
Figure 5: Comparison of test weighted-MAE on a) U-Net architecture and b) CNN autoencoder, in log scale, between the ground truth and prediction from the DNN surrogates across different acquisition functions as data acquisition proceeds.
Figure 6: Comparison of test weighted-MAE on a) U-Net architecture and b) CNN autoencoder, in log scale, between the ground truth and prediction from the DNN surrogates between data acquisitions using TOD and labelled-loss as data acquisition proceeds.
5.1 Comparison of Active Learning Acquisition Functions
Figure 6(a) compares weighted-MAE between the ground truth and DNN surrogate predictions for different acquisition functions at different percentages of labeled data, evaluated against random acquisitions for the U-Net architecture. As shown, actively selecting simulation data for DNN surrogates training, especially by considering the estimated loss of the surrogate (TOD), consistently improves the DNN performance compared to random acquisition using the same amount of labeled training data. Note that TOD, even though reliant on an estimate of the DNN loss without the simulation output as labels, acquires a performance on par with using the actual DNN loss calculated from labels as shown in Figure 6(a). In comparison, diversity-based selection of training samples demonstrated limited benefits with a performance similar to or marginally better than random acquisition.
Table  1 lists the weighted MAE between the simulated ground truth and predictions on the test data for U-Net across quarterly percentages of labeled data, computed for the different regions of interest as described in Figure 4. The results show that, at the same amount of labeled training data throughout the active learning process, the careful selection of training simulations achieved a test performance improvement across all acquisition strategies (except when all data are used): this improvement was more evident when the size of the labeled data was small. Additionally, these improvements were more pronounced at the regions proximal to the source (SRC, RING1) where the errors were the highest. This observation can be better appreciated in visual examples of the absolute errors (Figure 7) between DNN surrogate predictions and ground-truth simulation outputs obtained on the U-Net architecture. As expected, across the three rows, there is a gradual decrease in error as the training data increases. Notably, however, TOD consistently exhibits lower errors compared to random training, especially in regions close to the sources.
Figure 7: Comparison of absolute error maps between DNN predictions and the ground truth using between random (top row), diversity (middle row), and TOD (bottom row) acquisition function at different percentage of labelled data on U-Net architecture for two example lattice (1) and (2)
Figure 8: (a) Initial condition, (b) steady-state solution (ground truth), prediction of (c) U-Net and (e) CNN autoencoder surrogate, and prediction error of (d) U-Net and (f) CNN autoencoder surrogate, trained on 100% data, for two examples of two source simulation data where sources of 5-pixel radius are placed randomly within a 100 x 100 lattice. The scale for the error color bar for UNET and CNN autoencoder are different.
Figure 9: Comparison of absolute error maps between DNN predictions and the ground truth using between random (top row), diversity (middle row), and TOD (bottom row) acquisition function at different percentage of labelled data on CNN autoencoder. The scale for the error color bar is different from U-Net in Figure 7